OpenStudy (anonymous):

Evaluate (22,000,000)(3 • 104). Write the answer in scientific notation.

OpenStudy (anonymous):

i am going to assume that this is \[22,000,000\times 3.\times 10^4\]

OpenStudy (anonymous):

yeah i guess that works

OpenStudy (anonymous):

write \[22,000,000=2.2\times 10^7\] then multiply \[2.2\times 10^7\times3\times 10^4\] \[2.2\times 3\times 10^{7+4}\] etc

OpenStudy (anonymous):

okay thank you!